OpenStudy (anonymous):

is 10/20 and 20/40 equivalent or not equivalent

OpenStudy (agent47):

equivalent, since both are a half

OpenStudy (helpmeplzz):

both are equivalent

OpenStudy (agent47):

10 = 1*10 20 = 2*10 20=1*20 40=2*20

OpenStudy (anonymous):

you are right because both are 1/2